Write a linear equation in two variables that is satisfied by (-3, 5). Are the equations independent or dependant?
Answer by venugopalramana(3286)

Write a linear equation in two variables that is satisfied by (-3, 5). Are the equations independent or dependant?
PLEASE TYPE YOUR QUESTION CORRECTLY..
Write a linear equation ..ONE ..OR...TWO...?in two variables that is satisfied by (-3, 5). Are the equations....PLURAL HERE THAT IS 2 EQNS? independent or dependant?...QUESTION DOES NOT ARISE IF THERE IS ONLY 1 EQN.
OK..TAKING 2 EQNS....
1.....X+Y=2.....-3+5=2....OK.....................I
2.....X-Y=-8....-3-5=-8...OK.............................II
THEY ARE INDEPENDENT AS WE ARE GETTING ONE UNIQUE ANSWER SET FROM 2 EQNS. IN 2 UNKNOWNS.